OpenStudy (anonymous):

The lengths of the sides of a right triangle are a and b, and the hypotenuse is c. Find the area of the triangle. b = 2 in.; c = 6 in. A = sq. in.

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (studygurl14):

First find the length of side a.

jimthompson5910 (jim_thompson5910):

Area of Triangle 

jimthompson5910 (jim_thompson5910):

sorry I meant right triangle

OpenStudy (studygurl14):

\(\Large\tt Use~the~Pythagorean~Theorem\) \(\large a^2+b^2=c^2\)

OpenStudy (anonymous):

a is 1?

OpenStudy (studygurl14):

b= 2, and c=6...soo.... \(\large a^2+(2)^2=(6)^2\rightarrow a^2+4=36...\)

OpenStudy (studygurl14):

so..what does "a" equal?

OpenStudy (anonymous):

16?

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (studygurl14):

um...no \(\large a^2=32\rightarrow a=\sqrt{32}\rightarrow a=\sqrt{16\times2}\rightarrow 4\sqrt{2}\)

OpenStudy (anonymous):

i meant 16^2

OpenStudy (anonymous):

thnx